OpenStudy (anonymous):

A firecracker shoots up from a hill 145 feet high with an initial speed of 80 feet per second. Using the formula H(t) = −16t^2 + vt + s, approximately how long will it take the firecracker to hit the ground?

OpenStudy (anonymous):

i'll just put down the same method.. but you'll need to plug in different values for v and s into your quadratic.. http://openstudy.com/study#/updates/50119db9e4b08ffc0ef3d552

OpenStudy (anonymous):

i understand the formula but how do you solve for t?

OpenStudy (anonymous):

do you know how to use the quadratic formula?

OpenStudy (anonymous):

not really

OpenStudy (anonymous):

yes i know what it is but not good using it

OpenStudy (anonymous):

the solution to the quadratic equation: \(\large ax^2+bx+c=0 \) is obtained by the quadratic formula: \(\large x=\frac{-b\pm\sqrt{b^2-4ac}}{2a} \)

OpenStudy (anonymous):

first you need to identify your a, b, c: a = b = c = can you give me those values?

OpenStudy (anonymous):

a is 16 b is 145 and c is 80

OpenStudy (anonymous):

no.. be careful... a=-16 but there's more....

OpenStudy (anonymous):

thought i put the - my bad

OpenStudy (anonymous):

in the equation H(t) = −16t^2 + vt + s, v represents your initial velocity s represents your initial height

OpenStudy (anonymous):

so v is 80 and s is 145

OpenStudy (anonymous):

yes.... so a = b = c = ???

OpenStudy (anonymous):

a is -16 b is 80 c is 145

OpenStudy (anonymous):

yes.... now all you need to do is to stick these values into the quadratic formula.... let your calculator do the rest....

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

\[\huge x=\frac{-(80)\pm\sqrt{80^2-4(-16)(145)}}{2(-16)} \]

OpenStudy (anonymous):

yw...:)

OpenStudy (anonymous):

i forgot... the quadratic formula will give you two solutions.... choose the one that makes sense....

OpenStudy (anonymous):

ok